if the range is 22 and you hve the following numbers, what does the missing number have to be: 87, 76, 97, 78, 84, 83?
1
Expert's answer
2012-03-30T06:57:03-0400
The least number from the set is 76 and the greatest one is 97. 97 - 76 = 1, so there are two possibilities for the missing number: 76-1 = 75 or 97+1 = 98.